40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5909     Entry time: Wed Nov 16 10:25:57 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

The  PRM lost damping about a day ago. It was restored.

ELOG V3.1.3-